2.00 points Outback Outfitters sells recreational equipment One of the associated with the stove total $135,300 per month companys products, a smilil camp stove, selis for $110 per unit. Varisable expenses are $77 per stove, and fxed expenses Required 1 Compute the companys break-even point in unit sales and in dolar sales Number of stoves Total sales dolars 2 년 re variable expenses per stove ncrease os a percentage of the selling price, wa resa na h gher or . kwer break-ven porn (Altme rat teixes ei enses won unchanged) Higher break-even point Lower break-even point
2. If the variable expenses per stove increase unchanged.) as a percentage of the selling price, willit result in a higher or a lower break-even point? (Assume that the fixed expenses remain Higher break-even point Lower break-oven port 3. At pr sent. the company is seling 12.000 stoves per month. The sales manager is connced hat a 10% redet ninthe sengprce would esi na 5%ncrease stoves. Prepare two contribution format income statements, one under present operating conditions, and one as operations would appear after the proposed changes monysesof Present Proposed 12,000 Stoves Total Per unit Per unit
Proposed Present 12,000 Stoves Total Per unit Total Per unit 4. Rafer to the data in (3) above. How many stoves would have to be sold at the new seling price to yiold a minimum net operating income of $75,000 per month? (Round your answer to the nearest whole number )
Answer #1

Answer 1.

Break-Even Point in Unit Sales = Fixed Costs / (Sales Price Per Unit - Variable Costs Per Unit)

= $135300/( $110 - $77 ) = 4100 units.

Break-Even Point in Dollar Sales = Fixed cost / contribution margin

Contribution margin = (110 - 77 ) ÷110 = 0.3

Therefore, Break even point in Dollar Sales =

135300 ÷ 0.3 = $451000.

Answer 2.

If variable expense per stove increases as a percentage of sales price then it will result in higher break even point because if variable expense increases then contribution will decreses which results in increment of break even point.

Answer 3.

Income statement under present condition.

Sales $1320000 ( 12000 × $110 )
Less : variable cost $(924000) ( 12000 × $77 )
Contribution $396000
Less : Fixed cost $( 135300 )
Profit $260700

Income statement after proposed changes.

Sales

$1485000 ($99 × 15000 ) ( 10% reduction in selling price of $110 amounts to $99 and 25% increase in 12000 units is 15000 units.)

Less : variable cost ($1155000) ( $77 × 15000 )
Contribution $330000
Less : Fixed cost ($135300)
Profit $194700

Answer 4.

To yield a net operating income of $75000 per month, with new selling price, number of stoves required to be sold are 9559 units.